OpenStudy (anonymous):

Help PleAse!!! Complete the two-column proof. given:x/6+2=15 prove:X=78

OpenStudy (anonymous):

a. Given b. Subtraction Property of Equality c. Division Property of Equality a. Given b. Addition Property of Equality c. Multiplication Property of Equality a. Given b. Addition Property of Equality c. Division Property of Equality a. Given b. Subtraction Property of Equality c. Multiplication Property of Equality

OpenStudy (anonymous):

a: Given b: Subtraction Property of Equality c: Multiplication prpperty of Equality \[x/6+2=15\] \[x/6=13\] \[(6)x/6=13(6)\] \[x=78\]

OpenStudy (anonymous):

i think there is no such property as subtraction